OpenStudy (anonymous):

t a certain college the ratio to men to women is 6 to 5. If there are 1500 men, how many women are there

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

6/5 = 1500/x (set it u as a proportion and cross multiply) (6)(x) = (1500)(5) 6x = 7500 x = 7500/6 x = 1250 check... 6/5 = 1500/1250 (6)(1250) = (5)(1500) 7500 = 7500 (correct) there are 1250 women

OpenStudy (anonymous):

so you divide 6 and 5 ? right

OpenStudy (anonymous):

no.....I set it up as a proportion. 6 men to 5 women = 1500 men to x women. I then cross multiplied to find x (the number of women). Once I found x, I then checked my answer and it proved to be true. Can you understand how I did it now ?

OpenStudy (anonymous):

thanks u for helping me

OpenStudy (anonymous):

your welcome....thats what I am here for :)